OpenStudy (anonymous):

What is the sum of the multiples of 6 between 6 and 999? Here's what I did: n = 996/6 = 166? d = 6 t1 = 6 Sn = n/2 [2 (t1) + (n-1)d]

OpenStudy (raden):

the series is 12 + 8 + 24 + 30 + ... + 996 so, the frist term is 12

OpenStudy (raden):

looks u were right :) then use the formula above

OpenStudy (anonymous):

I did but it didn't work. The correct answer is 82, 665 but my answer was 83, 166.

OpenStudy (anonymous):

isn't the series supposed to be 6, 12, 18.....?

OpenStudy (anonymous):

82.665

OpenStudy (raden):

btw, there is a bit typo from me the series should is 12 + 18 + 24 + 30 + ... + 996 T1 = 12 (because the word "between 6 and 999" means that numbers are more than 6 and less than 999) d = 6 to get the value of n, u can find out by using the formula : Tn = T1 + (n-1)d 996 = 12 + (n-1)6 996 = 12 + 6n - 6 996 = 6n + 6 996 - 6 = 6n 990 = 6n n = 990/6 = 165 so, there are 165 terms thus, the sum all terms becomes Sn = n/2 [2 (t1) + (n-1)d] S165 =165/2 (2*12 + (165-1)6) S165 =165/2 (24 + 164*6) S165 =165/2 (24 + 984) S165 =165/2 (1008) S165 =83,160

OpenStudy (anonymous):

Let us put a Little logic in things. The solution cannot be 82.665 as a sum of multiples of 6 is also a multiple of 6 and an odd number cannot be multiple of 6. Therefore, the solution 82.665 is discarded. You have got the right solution :83.166

OpenStudy (raden):

@CarlosGP , do u know the means of "between 6 and 999" ? actually, it is